Six women and five men walk into an


appliance store at the same time. There are five


salesperson available to help them.Find the


probability that a salesperson will first help at


least 2 women.

1
Expert's answer
2022-02-24T17:08:16-0500

we define the total number of men and women = (5+6) = 11

we define (11C5) = 462

let X be the number of women to be helped first

p(x≥2) = 1 - ( P(x=0) + P(x=1) )

we define p(x=0) = ( (6C0) * (5C5) ) / (11C5) = 1/462

we define p(x=1) =( (6C1) * ( 5C4) ) / (11C5) = 30/462

Thus p(x≥2) = 1 - ( (1/462) + (30/462) )

= (431/462) = 0.932900432